### v1

# The maximal correlation coefficient associated with the minimum

Abstract
For independent random variables $(X_{i})_{1\leq i\leq n}$, we consider the maximal correlation coefficient $R=R(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j})$. If $X_{1},X_{2},\ldots,X_{n}$ are identically distributed with the same continuous distribution, we find that $R=(m-\ell)/\sqrt{m(n-\ell)}$. For discrete distributions, we calculate the maximal correlation coefficient $R$ for Bernoulli distributions, geometric distributions, binomial distributions and Poisson distributions. Our paper answers a question in [CC24, Section 6].
1 Introduction
The maximal correlation coefficient introduced by Gebelein in [Geb41] measures the dependency of two random variables. For two non-degenerate random variables $X$ and $Y$, it is given by
$R(X,Y)=\sup\frac{\operatorname{Cov}(\varphi(X),\psi(Y))}{\sqrt{\operatorname{Var}(\varphi(X))\operatorname{Var}(\psi(Y))}},$
where the supremum is taken over all measurable functions $\varphi$ and $\psi$ such that $\operatorname{Var}(\varphi(X)),\operatorname{Var}(\psi(Y))\in(0,\infty)$.
For the literature on the maximal correlation coefficient, we refer to our recent paper [CC24] and the references therein.
Among various results, Bücher and Staud [BS25, Corollary 2.2] have calculated the maximal correlation coefficient for the bivariate Marshall-Olkin exponential distribution. To be more precise, consider independent exponential random variables $X_{1},X_{2}$ and $X_{3}$ with parameters $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$, respectively. Then
$R(\min(X_{1},X_{2}),\min(X_{2},X_{3}))=\frac{\lambda_{2}}{\sqrt{(\lambda_{1}+\lambda_{2})(\lambda_{2}+\lambda_{3})}}.$
In [CC24, Section 5.6], we reproduce the upper bound
$R(\min(X_{1},X_{2}),\min(X_{2},X_{3}))\leq\frac{\lambda_{2}}{\sqrt{(\lambda_{1}+\lambda_{2})(\lambda_{2}+\lambda_{3})}}$ (1)
of their work, which follows as a consequence of the following proposition.
Proposition 1.1 (Proposition 5.5 in [CC24]).
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. (real) random variables. Let $1\leq\ell+1\leq m\leq n$. Then, we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)\leq\frac{m-\ell}{\sqrt{m(n-\ell)}}.$ (2)
By [BS25, Corollary 2.2], the upper bounds in (1) and (2) are sharp when each $X_{i}$ follows an exponential distribution. We thus wonder whether they are also sharp for other distributions, see the second open problem in [CC24, Section 6].
In the present paper, we prove that (2) is also sharp for other continuous distributions. However, it is not sharp for some discrete distributions. Moreover, we obtain explicit formulae for Bernoulli, geometric, binomial and Poisson distributions.
Theorem 1.2.
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. (real) continuous random variables. Let $1\leq\ell+1\leq m\leq n$. Then we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)=\frac{m-\ell}{\sqrt{m(n-\ell)}}.$
Theorem 1.3.
Let $X_{1},X_{2},\ldots,X_{n}$ be independent Bernoulli random variables such that $P(X_{i}=1)=1-P(X_{i}=0)=p_{i}\in(0,1)$ for $1\leq i\leq n$. Let $R=R\left(\min\limits_{i:1\leq i\leq m}X_{i},\min\limits_{j:\ell+1\leq j\leq n}X_{j}\right)$, where $1\leq\ell+1\leq m\leq n$. Then
$R=\frac{\prod_{i=1}^{n}p_{i}-\prod_{i=1}^{m}p_{i}\prod_{i=\ell+1}^{n}p_{i}}{\sqrt{\prod_{i=1}^{m}p_{i}(1-\prod_{i=1}^{m}p_{i})\prod_{i=\ell+1}^{n}p_{i}(1-\prod_{i=\ell+1}^{n}p_{i})}}.$ (3)
In particular, if $p_{1}=p_{2}=\cdots=p_{n}=p\in(0,1)$, we have $R=R_{m,\ell}(p)$, where
$R_{m,\ell}(p):=\frac{p^{n}-p^{m+n-\ell}}{\sqrt{p^{m+n-\ell}(1-p^{m})(1-p^{n-\ell})}}.$ (4)
Theorem 1.4.
Let $X_{1},X_{2},\ldots,X_{n}$ be independent geometric random variables. For $1\leq i\leq n$, the parameter of $X_{i}$ is $p_{i}\in(0,1)$. Let $1\leq\ell+1\leq m\leq n$. Then we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)$
$=\frac{\prod_{i=1}^{n}(1-p_{i})-\prod_{i=1}^{m}(1-p_{i})\prod_{i=\ell+1}^{n}(1-p_{i})}{\sqrt{\prod_{i=1}^{m}(1-p_{i})(1-\prod_{i=1}^{m}(1-p_{i}))\prod_{i=\ell+1}^{n}(1-p_{i})(1-\prod_{i=\ell+1}^{n}(1-p_{i}))}}.$
In particular, for $n=3$, $m=2$ and $\ell=1$, we have that
$R(\min(X_{1},X_{2}),\min(X_{2},X_{3}))$
$=\frac{\sqrt{(1-p_{1})(1-p_{3})}p_{2}}{\sqrt{(1-(1-p_{1})(1-p_{2}))(1-(1-p_{2})(1-p_{3}))}}.$
Theorem 1.5.
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. binomial random variables with parameter $(d,p)$ with $0<p<1$. Let $1\leq\ell+1\leq m\leq n$. Then we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)=R_{m,\ell}(1-(1-p)^{d}),$ (5)
where $R_{m,\ell}(p)$ is defined in (4).
Theorem 1.6.
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. Poisson random variables with parameter $\lambda>0$. Let $1\leq\ell+1\leq m\leq n$. Then we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)=R_{m,\ell}(1-e^{-\lambda}),$ (6)
where $R_{m,\ell}(p)$ is defined in (4).
Organization of the paper: In Section 2, we present three useful properties of the maximal correlation coefficient. In Section 3, we prove Theorem 1.2. In Section 4, we first prove Theorem 1.3. Then we prove Theorems 1.4 and 1.5 by using Theorem 1.3 and prove Theorem 1.6 by using Theorems 1.3 and 1.5.
2 Preliminary
We present three useful properties of the maximal correlation coefficient.
If $U=\varphi(X)$ is a measurable function of $X$ and $V=\psi(Y)$ is a measurable function of $Y$, then $R(U,V)\leq R(X,Y)$.
Suppose $(X_{k},Y_{k})$ converges to $(X,Y)$ in law. By [CC24, Lemma 2.3],
$R(X,Y)\leq\liminf_{k\to\infty}R(X_{k},Y_{k}).$ (7)
Lemma 2.1.
Consider a finite or countably infinite collection $(X_{i},Y_{i})_{i\in I}$ of independent random tuples. Then we have that
$R((X_{i})_{i\in I},(Y_{i})_{i\in I})=\sup_{i\in I}R(X_{i},Y_{i}).$
If the collection is finite, it is known as the Csáki-Fischer identity; see [CF63, Theorem 6.2] and [Wit75, Theorem 1]. For the case $|I|=\infty$, similar to the proof of [CC24, Theorem 1.1], we can prove Lemma 2.1 as a consequence of the Csáki-Fischer identity and [CC24, Lemma 2.3].
3 Continuous distribution
By [BS25, Corollary 2.2], the upper bound in (2) is sharp when each $X_{i}$ follows an exponential distribution. For general continuous random variables $X_{1},X_{2},\ldots,X_{n}$ with the common distribution function $F(x)$, define
$Y_{i}=\varphi(X_{i}),$
where $\varphi(x)=-\ln(1-F(x))$. Note that $\varphi$ is a non-decreasing function. Moreover, $Y_{1},Y_{2},\ldots,Y_{n}$ are i.i.d. exponential random variables with parameter 11. Hence, we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)$ $\geq R\left(\varphi\left(\min_{i:1\leq i\leq m}X_{i}\right),\varphi\left(\min_{j:\ell+1\leq j\leq n}X_{j}\right)\right)$
$=R\left(\min_{i:1\leq i\leq m}\varphi(X_{i}),\min_{j:\ell+1\leq j\leq n}\varphi(X_{j})\right)$
$=R\left(\min_{i:1\leq i\leq m}Y_{i},\min_{j:\ell+1\leq j\leq n}Y_{j}\right)$
$=\frac{m-\ell}{\sqrt{m(n-\ell)}}.$ (8)
The proof is complete by combining the lower bound (3) with the upper bound (2).
Remark 3.1.
Alternatively, the lower bound (3) could be derived from Theorem 1.3. Indeed, let
$h(x)=1_{(\delta(p),\infty)}(x),$
where $P(X_{i}>\delta(p))=p$. Then
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)$ $\geq R\left(h\left(\min_{i:1\leq i\leq m}X_{i}\right),h\left(\min_{j:\ell+1\leq j\leq n}X_{j}\right)\right)$
$=R\left(\min_{i:1\leq i\leq m}h(X_{i}),\min_{j:\ell+1\leq j\leq n}h(X_{j})\right)$
[…]
$\overset{p\to 1}{\to}\frac{m-\ell}{\sqrt{m(n-\ell)}}.$
4 Some special discrete distributions
4.1 Bernoulli distribution
In this subsection, we prove Theorem 1.3.
Firstly, we calculate the maximal correlation coefficient between two Bernoulli random variables. The following lemma is taken from [CC24, Section 3.2].
Lemma 4.1.
Consider a random vector $\left(\xi,\eta\right)$ with the joint distribution in Table 1.
$\xi\backslash\eta$ $c$ $d$
$a$ $p_{ac}$ $p_{ad}$
$b$ $p_{bc}$ $p_{bd}$
Table 1: The joint distribution of $(\xi,\eta)$
Then we have that
$R(\xi,\eta)=\left.\left|\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}\right.\right|/\sqrt{p_{a}p_{b}p_{c}p_{d}},$
where $p_{a}=P(\xi=a)$, $p_{b}=P(\xi=b)$, $p_{c}=P(\eta=c)$ and $p_{d}=P(\eta=d)$.
Next, consider independent Bernoulli random variables $X_{1},X_{2},\ldots,X_{n}$, such that $P(X_{i}=1)=1-P(X_{i}=0)=p_{i}$ for $1\leq i\leq n$. Let $X=\min_{i:1\leq i\leq m}X_{i}$, $Y=\min_{j:\ell+1\leq j\leq n}X_{j}$. Then $X$ and $Y$ are both Bernoulli random variables with the joint distribution described in Table 2.
[…]
0 $1-{\textstyle\prod_{i=\ell+1}^{n}}p_{i}-{\textstyle\prod_{i=1}^{m}}p_{i}+{\textstyle\prod_{i=1}^{n}}p_{i}$ ${\textstyle\prod_{i=\ell+1}^{n}}p_{i}-{\textstyle\prod_{i=1}^{n}}p_{i}$
11 ${\textstyle\prod_{i=1}^{m}}p_{i}-{\textstyle\prod_{i=1}^{n}}p_{i}$ ${\textstyle\prod_{i=1}^{n}}p_{i}$
Table 2: The joint distribution of $(X,Y)$
By Lemma 4.1, we have that
$R(X,Y)=\frac{\prod_{i=1}^{n}p_{i}-\prod_{i=1}^{m}p_{i}\prod_{i=\ell+1}^{n}p_{i}}{\sqrt{\prod_{i=1}^{m}p_{i}(1-\prod_{i=1}^{m}p_{i})\prod_{i=\ell+1}^{n}p_{i}(1-\prod_{i=\ell+1}^{n}p_{i})}}.$
Remark 4.1.
Recall the definition of $R_{m,\ell}(p)$ in (4). Note that $R_{m,\ell}(p)$ is continuous on $(0,1)$ with $R_{m,\ell}(0+)=0$ and $R_{m,\ell}(1-)=(m-\ell)/\sqrt{m(n-\ell)}$. Moreover, $R_{m,\ell}(p)$ is non-decreasing.
Proof.
From (4), it is not difficult to check the continuity of $p\mapsto R_{m,\ell}(p)$ and to obtain the values of $R_{m,\ell}(0+)$ and $R_{m,\ell}(1-)$. For the monotonicity of $R_{m,\ell}$, let $a=(n+\ell-m)/2$, $b=m$-$\ell$ and $c=n$-$\ell$. Then for $0<p<1$,
$\frac{R_{m,\ell}^{\prime}(p)}{R_{m,\ell}(p)}$ $=\frac{1}{p}\left(a-\frac{bp^{b}}{1-p^{b}}+\frac{mp^{m}}{2(1-p^{m})}+\frac{cp^{c}}{2(1-p^{c})}\right)$
$\overset{2a+2b-m-c=0}{=}\frac{1}{2p}\left(\left(\frac{m}{1-p^{m}}-\frac{b}{1-p^{b}}\right)+\left(\frac{c}{1-p^{c}}-\frac{b}{1-p^{b}}\right)\right).$
For $x>0$, define $f(x)=x/(1-p^{x})$. Then
$f^{\prime}(x)=\frac{1-p^{x}+xp^{x}\ln p}{(1-p^{x})^{2}}>0.$
Hence, $f(x)$ is increasing. Since $b\leq m$, we have that $f(m)-f(b)\geq 0$. Similarly, $f(c)-f(b)\geq 0$. Therefore, we have that
$\frac{R_{m,\ell}^{\prime}(p)}{R_{m,\ell}(p)}=\frac{1}{2p}(f(m)-f(b)+f(c)-f(b))\geq 0,$
and $R_{m,\ell}(p)$ is non-decreasing. ∎
4.2 Geometric distribution
In this subsection, we prove Theorem 1.4. Since the geometric distribution and the Bernoulli distribution are closely related, we will deduce Theorem 1.4 from Theorem 1.3.
Consider independent Bernoulli random variables $(B_{ik})_{i,k\geq 1}$ such that $B_{ik}$ has the parameter $1-p_{i}$ for $i,k\geq 1$. For $i\geq 1$, let $X_{i}=\min\{k\geq 1:B_{ik}=0\}$. Then $(X_{i})_{i\geq 1}$ is a sequence of independent geometric random variables such that $X_{i}$ has the parameter $p_{i}$ for $i\geq 1$.
Let $U=\min_{i:1\leq i\leq m}X_{i}$ and $V=\min_{j:\ell+1\leq j\leq n}X_{j}$. Then
$U=\min\{k\geq 1:U_{k}=0\},$
where
$U_{k}=\min_{i:1\leq i\leq m}B_{ik}.$
Similarly,
$V=\min\{k\geq 1:V_{k}=0\},$
where
$V_{k}=\min_{j:\ell+1\leq j\leq n}B_{jk}.$
On the other hand,
$U_{1}=1_{U\geq 2}\text{ and }V_{1}=1_{V\geq 2}.$
By Theorem 1.3, for $k\geq 1$, we have that
$R(U_{k},V_{k})=R(U_{1},V_{1})$
$=\frac{\prod_{i=1}^{n}(1-p_{i})-\prod_{i=1}^{m}(1-p_{i})\prod_{i=\ell+1}^{n}(1-p_{i})}{\sqrt{\prod_{i=1}^{m}(1-p_{i})(1-\prod_{i=1}^{m}(1-p_{i}))\prod_{i=\ell+1}^{n}(1-p_{i})(1-\prod_{i=\ell+1}^{n}(1-p_{i}))}}.$
Since $(U_{k},V_{k})_{k\geq 1}$ is a sequence of i.i.d. random vectors, by Lemma 2.1, we have that
$R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1})=\sup_{k\geq 1}R(U_{k},V_{k})=R(U_{1},V_{1}).$
Finally, since $U$ is a measurable function of $(U_{k})_{k\geq 1}$ and $V$ is a measurable function of $(V_{k})_{k\geq 1}$, we have that
$R(U,V)\leq R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1}).$
On the other hand, since $U_{1}$ is a measurable function of $U$ and $V_{1}$ is a measurable function of $V$, we have that
$R(U,V)\geq R(U_{1},V_{1}).$
Finally, we conclude that
$R(U,V)=R(U_{1},V_{1})$
$=\frac{\prod_{i=1}^{n}(1-p_{i})-\prod_{i=1}^{m}(1-p_{i})\prod_{i=\ell+1}^{n}(1-p_{i})}{\sqrt{\prod_{i=1}^{m}(1-p_{i})(1-\prod_{i=1}^{m}(1-p_{i}))\prod_{i=\ell+1}^{n}(1-p_{i})(1-\prod_{i=\ell+1}^{n}(1-p_{i}))}}.$
4.3 Binomial distribution
In this subsection, we prove Theorem 1.5 by using Theorem 1.3.
Consider i.i.d. binomial random variables $X_{1},X_{2},\ldots,X_{n}$ with parameter $(d,p)$. Let $(B_{ik})_{i,k\geq 1}$ be independent Bernoulli random variables such that
$P(B_{ik}=1)=p(k):=P(X_{i}\geq k|X_{i}\geq k-1).$
We need the following lemma on the monotonicity of $k\mapsto p(k)$.
Lemma 4.2.
The conditional probability $p(k)$ is non-increasing in $k$. In particular, for all $k\geq 1$, $p(k)\leq p(1)=1-(1-p)^{d}$.
Proof.
Note that
$\frac{p(k+1)}{p(k)}=\frac{P(X_{i}\geq k+1)P(X_{i}\geq k-1)}{\left(P(X_{i}\geq k)\right)^{2}}.$
Hence, it suffices to prove that
$\Delta:=\left(P(X_{i}\geq k)\right)^{2}-P(X_{i}\geq k+1)P(X_{i}\geq k-1)\geq 0$
for $k\geq 1$. Note that
$P(X_{i}\geq k)=P(X_{i}=k)+P(X_{i}\geq k+1),$
$P(X_{i}\geq k-1)=P(X_{i}=k-1)+P(X_{i}\geq k).$
Then we have that
$\Delta=P(X_{i}=k)P(X_{i}\geq k)-P(X_{i}=k-1)P(X_{i}\geq k+1).$
For $k\geq 1$ and $j\geq 0$, since
$\binom{n}{k}=\frac{n-k+1}{k}\binom{n}{k-1}\text{ and }\binom{n}{k+j+1}=\frac{n-k-j}{k+j+1}\binom{n}{k+j}$,
we have that
$\binom{n}{k}\binom{n}{k+j}\geq\binom{n}{k-1}\binom{n}{k+j+1}$.
Therefore, we have that
$P(X_{i}=k)P(X_{i}=k+j)\geq P(X_{i}=k-1)P(X_{i}=k+j+1).$
By summing over $j\geq 0$, we see that $\Delta\geq 0$, and the proof is complete. ∎
Let us continue with the proof of Theorem 1.5. We have the following equality in distribution:
$X_{i}\overset{(d)}{=}\min\{k\geq 1:B_{ik}=0\}-1.$
Let $U=\min_{i:1\leq i\leq m}X_{i}$ and $V=\min_{j:\ell+1\leq j\leq n}X_{j}$. Then
$U=\min\{k\geq 1:U_{k}=0\}-1,$
where
$U_{k}=\min_{i:1\leq i\leq m}B_{ik}.$
Similarly,
$V=\min\{k\geq 1:V_{k}=0\}-1,$
where
$V_{k}=\min_{j:\ell+1\leq j\leq n}B_{jk}.$
On the other hand,
$U_{1}=1_{U>0}\text{ and }V_{1}=1_{V>0}.$
Since $U_{1}$ is a measurable function of $U$ and $V_{1}$ is a measurable function of $V$, we have that
$R(U_{1},V_{1})\leq R(U,V).$
Since $U$ is a measurable function of $(U_{k})_{k\geq 1}$ and $V$ is a measurable function of $(V_{k})_{k\geq 1}$, we have that
$R(U,V)\leq R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1}).$
By Lemma 2.1, Theorem 1.3, Remark 4.1 and Lemma 4.2,
$R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1})=\sup_{k\geq 1}R(U_{k},V_{k})=\sup_{k\geq 1}R_{m,\ell}(p(k))=R_{m,\ell}(1-(1-p)^{d}).$
4.4 Poisson distribution
In this subsection, we deduce Theorem 1.6 from Theorems 1.3 and 1.5.
For $k\geq 1$, let $X^{(k)}_{1},X^{(k)}_{2},\ldots,X^{(k)}_{n}$ be i.i.d. binomial random variables with parameter $(k,\lambda/k)$. Then $(X^{(k)}_{1},X^{(k)}_{2},\ldots,X^{(k)}_{n})$ converges in distribution towards $(X_{1},X_{2},\ldots,X_{n})$ as $k\to\infty$, where $X_{1},X_{2},\ldots,X_{n}$ are i.i.d. Poisson random variables with parameter $\lambda$. By Theorem 1.5, we have that
$R\left(\min_{i:1\leq i\leq m}X^{(k)}_{i},\min_{j:\ell+1\leq j\leq n}X^{(k)}_{j}\right)=R_{m,\ell}(1-(1-\lambda/k)^{k}).$
By taking $k\to\infty$, using (7) and the continuity of $R_{m,\ell}$, we obtain that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)\leq R_{m,\ell}(1-e^{-\lambda}).$
On the other hand, we take $Y_{i}=1_{X_{i}>0}$. Then $Y_{1},Y_{2},\ldots,Y_{n}$ are i.i.d. Bernoulli random variables with parameter $1-e^{-\lambda}$. Moreover, the Bernoulli random variable $\min_{i:1\leq i\leq m}Y_{i}$ is equal to 0 if and only if $\min_{i:1\leq i\leq m}X_{i}=0$. Hence, $\min_{i:1\leq i\leq m}Y_{i}$ is a measurable function of $\min_{i:1\leq i\leq m}X_{i}$. Similarly, the random variable $\min_{j:\ell+1\leq j\leq n}Y_{j}$ is a measurable function of $\min_{j:\ell+1\leq j\leq n}X_{j}$. Hence, by Theorem 1.3, we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)\geq R\left(\min_{i:1\leq i\leq m}Y_{i},\min_{j:\ell+1\leq j\leq n}Y_{j}\right)=R_{m,\ell}(1-e^{-\lambda}).$